Tickets for a show cost R 500 or R 800. If 100 tickets were sold for a total of R 65000, how many
tickets were sold for 800?

Answer by JulietG(1812)

If all 100 tickets were sold at R800, the sales total would be 100 * 800 or R80000.
The actual total is R65000. R80000 - R65000 = R15000
The difference in price between R800 and R500 is R300.
R15000 / R300 = 50 tickets
50 tickets at R500 = R25000
50 tickets at R800 = R40000
R25000 + R40000 = R65,000